OpenStudy (anonymous):

Assume the random variable X is normally distributed with u=50 and standard deviation o=7. compute the probability. P(X>34)=? please explain how to get the answer.

OpenStudy (phi):

to do your problem we need to first find the z-score. the video explains that part.

OpenStudy (anonymous):

well from what i got from the video i think i do \[\frac{ 34-50 }{ 7 } = \frac{ -16 }{ 7 } = -2.29\]

OpenStudy (phi):

yes. now we need to change that into a probability. I am looking at his videos trying to find the one that explains how to do it.

OpenStudy (anonymous):

well using this for - z scores i got 0.011011

OpenStudy (phi):

yes, so you know this part! you get about 1.1% chance or 0.011

OpenStudy (anonymous):

lol yah & thx!

OpenStudy (anonymous):

what answer do i put? i put 0.0110 but it says i'm wrong

OpenStudy (anonymous):

oh!! i think i got it

OpenStudy (phi):

oops, I wasn't thinking. we found the area under the curve from -infinity up to 34. in other words we found Pr(x≤34)